# Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism

### v1

@@ Abstract
Quantum Chromodynamics (QCD) admits a topological $\theta$-term that violates Charge-Parity (CP) symmetry, yet experimental indicate that $\theta$ is nearly zero. To investigate this discrepancy in a controlled setting, we derive the Hamiltonian representation of the QCD Lagrangian and construct its (1+1)(1{+}1)-dimensional Schwinger-model analogue. By encoding fermionic and gauge degrees of freedom into qubits using the Jordan-Wigner and quantum-link schemes, we obtain a compact Pauli Hamiltonian that retains the relevant topological vacuum structure. Ground states are prepared using a feedback-based quantum optimization protocol, enabling numerical evaluation of the vacuum energy $E_{0}(\theta)$ on a few-qubit simulator. Our results show a displaced vacuum at nonzero $\theta$ in agreement with strong-interaction expectations, and demonstrate that introducing a dynamical axion field drives the system toward $\theta=0$, thereby realizing the Peccei-Quinn mechanism within a minimal quantum simulation. These results illustrate how quantum hardware can examine symmetry violation and its dynamical resolution in gauge theories.
@@
Introduction. Quantum Chromodynamics (QCD) describes the fundamental interactions of quarks and gluons through the non-Abelian SU(3) gauge symmetry [https://doi.org/10.1002/andp.200051211-1210, Gross2023, doi:10.1142/S0217732324501943]. The QCD Lagrangian captures both the dynamics of quark fields and the gluon self-interactions that arise from this non-Abelian structure [doi:10.1142/S0217732324501943]. In addition to these standard kinetic and interaction terms, QCD also permits a topological contribution that explicitly violates Charge Parity ($CP$) symmetry [PhysRevLett.37.8, PhysRevD.14.3432]. Although this term does not affect perturbative processes, it plays a central role in shaping the nonperturbative vacuum structure of the theory [RevModPhys.82.557]. The most general QCD Lagrangian, including the $CP$-violating topological term, can be written as
$\mathcal{L}_{\text{QCD}}=-\frac{1}{4}G_{\mu\nu}^{a}G^{a,\mu\nu}+\sum_{f}$ $\bar{\psi}_{f}(i\gamma^{\mu}D_{\mu}-m_{f})\psi_{f}$
$+\theta\frac{g^{2}}{32\pi^{2}}G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu},$ (1)
@@
where $\psi_{f}$ denotes the quark field of flavor $f$ with mass $m_{f}$, $g$ is the strong coupling constant, and $\tilde{G}^{a,\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}G_{\rho\sigma}^{a}.$ The gluon field strength tensor is defined by $G_{\mu\nu}^{a}=\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{a}+gf^{abc}A_{\mu}^{b}A_{\nu}^{c},$ where $A_{\mu}^{a}$ are the gluon gauge fields and $f^{abc}$ are the structure constants of the SU(3) Lie algebra. The covariant derivative acting on the quark fields takes the form $D_{\mu}=\partial_{\mu}-igT^{a}A_{\mu}^{a},$ with $T^{a}$ representing the SU(3) generators satisfying the commutation relation $T^{a},T^{b}]=if^{abc}T^{c}$.
The first term in Eq. (Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism) corresponds to the kinetic and self-interaction energy of gluons, reflecting the inherent nonlinearity of non-Abelian gauge theories. The second term describes the dynamics of quarks, their Dirac mass contributions, and their interactions with gluons mediated by the color charge. The final term, proportional to the dimensionless parameter $\theta$, represents the strong $CP$-violating term. Unlike the electroweak theory, QCD allows a topological contribution that originates from multi-sector vacuum. In the path-integral picture, the vacuum is a sum over gauge configurations labeled by a winding number $n$ connect these sectors. The $\theta$-term contributes a phase $\exp(in\theta)$ to each topological sector, which shifts the vacuum toward $CP$-violating configurations [PhysRevLett.37.8, PhysRevLett.37.172, PhysRevD.86.105020].
@@
The tensor structure $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}$ is odd under parity $P$ and charge-parity $CP$ transformations, therefore any nonzero $\theta$ introduces explicit $CP$ violation in the strong interaction. In QCD, such a violation would be expected to generate low-energy hadronic effects; for example, it induces a permanent electric dipole moment (EDM) of the neutron [CREWTHER1979123, POSPELOV2005119]. Precision searches, however, have not observed such a signal. The current experimental bound places the neutron EDM $|d_{n}|<1.8\times 10^{-26}e\cdot\mathrm{cm}\ (90\%{\rm C.L.})$ [PhysRevLett.97.131801, PhysRevLett.124.081803], which implies an extraordinarily small value of $|\theta|\lesssim 10^{-10}$. This extreme suppression cannot be explained within the Standard Model, leading to the strong $CP$ problem: QCD allows a large $\theta$, yet nature appears to choose a value extremely close to zero.
@@
A compelling theoretical resolution is provided by the Peccei-Quinn mechanism [PhysRevLett.38.1440, PhysRevD.16.1791, Peccei2008, RevModPhys.82.557], which promotes $\theta$ to a dynamical field associated with a spontaneously broken global $U(1)_{\text{PQ}}$ symmetry. The resulting pseudo-Nambu-Goldstone boson, known as the axion, dynamically relaxes the effective $\theta$ parameter to zero, thereby restoring $CP$ symmetry in the strong interactions [RevModPhys.82.557]. The axion framework not only resolves the strong $CP$ problem but also provides a well-motivated candidate for dark matter, linking fundamental particle physics with cosmological observations [doi:10.1126/sciadv.abj3618].
@@
Theoretical progress on the strong-$CP$ problem is limited by the fact that the QCD vacuum is intrinsically nonperturbative and difficult to treat with conventional techniques. Conventional lattice methods struggle with real-time dynamics and $\theta$-dependence, while analytical techniques rely on approximations. Quantum simulation provides a complementary route: instead of performing numerical sampling, one implements a controllable quantum system whose Hilbert space directly encodes the gauge theory and then measures how its vacuum responds to a $CP$-violating background. The (1+1)(1{+}1)-dimensional Schwinger model retains essential features such as confinement and topological vacuum sectors, yet is sufficiently simple to be mapped to a small number of qubits and executed on near-term hardware.
@@
In this work, we use this model to examine the microscopic origin of strong $CP$ violation and its axionic restoration. We construct a qubit-encoded Schwinger model with a $\theta$-dependent electric sector, extend it to include a dynamical axion field, and prepare its vacuum using a feedback-driven quantum optimization algorithm (FALQON) [PhysRevLett.129.250502]. By computing the ground-state energy as a function of $\theta$, we first identify the displaced vacuum expected for $CP$-violating gauge theories and then demonstrate its relaxation toward $\theta=0$ once the axion is activated. These results show that the Peccei-Quinn mechanism, the dynamical cancellation of $CP$ violation, can be realized in controlled, few-qubit settings, suggesting a new avenue for probing topological gauge physics with quantum simulation.
@@
Lagrangian to Hamiltonian. To express the QCD Lagrangian in a form suitable for canonical quantization, we rewrite the gluon field-strength tensor $G_{\mu\nu}^{a}$ in terms of its temporal and spatial components. In analogy to electromagnetism, we define the chromoelectric and chromomagnetic fields as
@@
Substituting these definitions into the gauge-field part of the QCD Lagrangian
$\mathcal{L}_{\text{gauge}}=-\frac{1}{4}G_{\mu\nu}^{a}G^{a,\mu\nu}+\theta\frac{g^{2}}{32\pi^{2}}G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu},$ (3)
and using the identity $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}=-4\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, we obtain the Lagrangian density expressed in terms of the physical fields
$\mathcal{L}_{\text{gauge}}=\frac{1}{2}\mathbf{E}^{a}\cdot\mathbf{E}^{a}-\frac{1}{2}\mathbf{B}^{a}\cdot\mathbf{B}^{a}+\theta\frac{g^{2}}{8\pi^{2}}\mathbf{E}^{a}\cdot\mathbf{B}^{a}.$ (4)
@@
$\Pi_{i}^{a}=\frac{\partial\mathcal{L}_{\text{gauge}}}{\partial(\partial_{0}A_{i}^{a})}=E_{i}^{a}+\theta\frac{g^{2}}{8\pi^{2}}B_{i}^{a},$ (5)
showing that the $\theta$-term shifts the momentum by an amount proportional to the chromomagnetic field. This term couples the electric and magnetic sectors, thus $\Pi_{i}^{a}$ no longer equals the electric field, and thereby modifying the canonical structure of the theory.
The Hamiltonian is then constructed via a Legendre transformation [10.1119/1.3119512]
@@
$=\frac{1}{2}\left(\mathbf{E}^{a}\cdot\mathbf{E}^{a}+\mathbf{B}^{a}\cdot\mathbf{B}^{a}\right)-\theta\frac{g^{2}}{8\pi^{2}}\mathbf{E}^{a}\cdot\mathbf{B}^{a}.$ (6)
The first two terms represent the standard gauge-field Hamiltonian corresponding to pure QCD without $CP$ violation. The final term, linear in the scalar product $\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, arises directly from the $\theta$-term and encapsulates its $CP$-odd character.
@@
Consequently, the term proportional to $\theta\mathbf{E}^{a}\cdot\mathbf{B}^{a}$ in the Hamiltonian explicitly breaks $CP$ symmetry. Although small in magnitude (given experimental constraints on $\theta$), this term could induce observable effects such as a permanent EDM in color-neutral hadrons like the neutron [CREWTHER1979123, POSPELOV2005119].
Simplified (1+1)D U(1) Schwinger model. We consider a (1+1)-dimensional Schwinger model, corresponding to one spatial and one temporal dimension [PhysRevD.52.6435, PhysRev.128.2425]. Despite its Abelian gauge structure, this model reproduces several nonperturbative features in QCD, including confinement of charges, a dynamically generated mass gap, and spontaneous chiral symmetry breaking in the massless limit [PhysRev.128.2425, Luo2007, PhysRevD.104.126029, PhysRevD.101.054507]. It also possesses a nontrivial topological vacuum structure analogous to the $\theta$-vacua of QCD, allowing controlled investigations of $CP$-violating effects induced by a background $\theta$-term in (1+1) dimensions [PhysRevD.101.054507, PhysRevD.46.5598]. Crucially, the Schwinger model admits a well-defined lattice discretization and a finite-dimensional gauge-link formulation [nt76-ttmj], making it ideally suited for implementation on quantum computing platforms.
In (1+1) dimensions, the gauge field has only one spatial component, and the magnetic field identically vanishes. The gauge-field dynamics are therefore governed by the electric field $E(x)$, which serves as the canonical conjugate to the spatial gauge potential $A_{1}(x)$. The commutation relation reads $A_{1}(x),E(y)]=i\delta(x-y)$, analogous to that between position and momentum in quantum mechanics. The continuum Lagrangian for a single fermion flavor coupled to this Abelian gauge field, including the topological $\theta$-term, can be written as
$\mathcal{L}=\bar{\psi}\big(i\gamma^{\mu}D_{\mu}-m\big)\psi-\frac{1}{2}E^{2}+\frac{\theta}{2\pi}E,$ (9)
where $D_{\mu}=\partial_{\mu}-igA_{\mu}$ denotes the gauge-covariant derivative and $g$ is the coupling constant. Here, we omit the subscript $f$ for short, such as $m\equiv m_{f}$. The first term describes the relativistic dynamics of the fermion field $\psi$, the second term represents the electric-field energy density, and the final term is the (1+1)D analog of the strong $CP$ term in QCD. This $\theta$-term corresponds physically to a uniform background electric field, generating $CP$-violating effects analogous to those of the $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}$ term in the non-Abelian case.
A Legendre transformation gives the Hamiltonian form
$H=\int dx\Big[\psi^{\dagger}(x)\big(-i\gamma^{1}D_{1}+m\gamma^{0}\big)\psi(x)+\tfrac{1}{2}\Big(E(x)-\tfrac{\theta}{2\pi}\Big)^{2}\Big],$
where the shift $E(x)\!\to\!E(x)-\tfrac{\theta}{2\pi}$ encodes the $CP$-odd topological term. This structure is analogous to the $\mathbf{E}\!\cdot\!\mathbf{B}$ coupling in QCD but simpler in one spatial dimension.
For quantum simulation, we adopt the Kogut–Susskind lattice discretization [PhysRevD.11.395, PhysRevD.91.054506], with gauge links $U_{x,x+1}=e^{iA_{x,x+1}}$ and electric fields $E_{x}$. The discretized Hamiltonian reads
@@ A Legendre transformation gives the Hamiltonian form
$+\frac{g^{2}}{2}\sum_{x}\Big(E_{x}-\frac{\theta}{2\pi}\Big)^{2},$ (10)
where $Z_{x}\!\equiv\!1-2\psi_{x}^{\dagger}\psi_{x}$ denotes the local charge operator. The alternating sign $(-1)^{x}$ stems from the staggered-fermion scheme, which suppresses fermion doubling while preserving chiral symmetry. The $\theta$-shift produces a uniform $CP$-violating background analogous to that of QCD.
@@
Two lattice sites case. Restricting the system to its smallest nontrivial instance, i.e., two sites lattice connected by a single link, yields the minimal Hamiltonian
@@
$+\dfrac{g^{2}}{2}\left(E-\frac{\theta}{2\pi}\right)^{2}+\lambda\sum_{x}G_{x}^{2}.$ (12)
In this representation, the first term corresponds to the staggered mass energy, the second encodes gauge-invariant fermion hopping between the two sites mediated by the link operator $U_{01}$, and the third term represents the electric-field energy including the $CP$-violating $\theta$ shift. The final term enforces Gauss’s law through an energetic constraint, restricting the evolution to the gauge-invariant subspace.
This minimal Schwinger model retains the topological and $CP$-violating structure of the QCD Hamiltonian while being computationally tractable on near-term quantum processors. The $\theta$-dependent electric-field shift plays the role of an effective $CP$-odd background field, allowing one to explore parity violation, vacuum degeneracy, and $\theta$-vacuum transitions within a low-dimensional analog of strong interactions.
To implement the model on a quantum computer, both the fermionic and gauge degrees of freedom are encoded in qubits (spin-$\tfrac{1}{2}$ systems). The fermionic modes are mapped to qubits using the Jordan–Wigner (JW) [Jordan1928], which preserves the correct anticommutation relations through parity strings, while each gauge link is represented by a single qubit in the quantum-link formulation. The staggered fermion fields $\psi_{x}$ and $\psi_{x}^{\dagger}$ satisfy
@@
where $S_{x}^{\pm}=(X_{x}\pm iY_{x})/2$. The string $\prod_{y<x}Z_{y}$ enforces fermionic parity between sites. Here $X$, $Y$, and $Z$ are Pauli matrices. The quantum-link representation of the gauge field is given by $E=(I-Z_{\ell})/2$ and $U=(X_{\ell}+iY_{\ell})/2$, where $\ell$ stands for the single-link qubit. See the End Matter section for the full derivation.
@@
The electric-field energy term, including the $CP$-violating $\theta$-shift yields
$H_{\theta}$ $=\dfrac{g^{2}}{2}\Big(E-\dfrac{\theta}{2\pi}\Big)^{2}$
$=\frac{g^{2}}{2}\left[\left(\frac{\theta}{2\pi}\right)^{2}-\frac{\theta}{2\pi}+\frac{1}{2}\right]I_{l}+\frac{g^{2}}{2}\left(\frac{\theta}{2\pi}-\frac{1}{2}\right)Z_{\ell}.$ (16)
The constant term merely shifts the global energy, while the $Z_{\ell}$-dependent term changes sign at $\theta=\pi$, leading to the correct $2\pi$ periodicity.
@@
The first term encodes the fermionic mass asymmetry, the second represents correlated fermion-gauge hopping, the third incorporates the $\theta$-dependent $CP$-odd energy shift, and the last enforces gauge invariance. Despite its simplicity, this model retains the nontrivial topological and $CP$-violating structure of QCD and provides an ideal platform for near-term quantum simulations of gauge-theory dynamics.

### v2

@@ Abstract
Quantum Chromodynamics (QCD) admits a topological $\bar{\theta}$-term that violates Charge-Parity $(CP)$ symmetry, yet experimental indicate that $\bar{\theta}$ is nearly zero. To investigate this discrepancy in a controlled setting, we derive the Hamiltonian representation of the QCD Lagrangian and construct its (1+1)(1{+}1)-dimensional Schwinger-model analogue. By encoding fermionic and gauge degrees of freedom into qubits using the Jordan-Wigner and quantum-link schemes, we obtain a compact Pauli Hamiltonian that retains the relevant topological vacuum structure. Ground states are evaluated using a feedback-based quantum optimization protocol, enabling numerical evaluation of the vacuum energy $E_{0}(\bar{\theta})$ on a few-qubit simulator. Our results show a vacuum at $\bar{\theta}=0\ (\text{mod 2})$ in agreement with strong-interaction expectations, and demonstrate that introducing a dynamical axion field drives the system toward a $\theta_{\rm eff}=0$, thereby realizing the Peccei-Quinn mechanism within a minimal quantum simulation. These results illustrate how quantum hardware can examine symmetry violation and its dynamical resolution in gauge theories.
@@
Introduction. Quantum chromodynamics (QCD) is an SU(3) non-Abelian gauge theory describing the interactions of quarks and gluons [1, 2, 3]. In addition to the standard kinetic and interaction terms [3], the QCD action admits a topological term associated with gauge-field configurations of nontrivial winding number. This term is odd under charge-parity $(CP)$ transformations and leads to explicit $CP$ violation when its coefficient is nonzero [4, 5]. Although this term does not affect perturbative processes, it plays a central role in shaping the nonperturbative vacuum structure of the theory [6]. Including this contribution, the QCD Lagrangian can be written as
$\mathcal{L}_{\rm QCD}=-\frac{1}{4}G_{\mu\nu}^{a}G^{a,\mu\nu}+\sum_{f}\bar{\psi}_{f}\left(i\gamma^{\mu}D_{\mu}-m_{f}\right)\psi_{f}$
$+\bar{\theta}\,\frac{g^{2}}{32\pi^{2}}\,G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}.$ (1)
@@
Here $G_{\mu\nu}^{a}=\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{a}+gf^{abc}A_{\mu}^{b}A_{\nu}^{c}$ denotes the gluon field-strength tensor, and $\tilde{G}^{a,\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}G_{\rho\sigma}^{a}$ its dual, where $A_{\mu}^{a}$ are the gluon gauge fields and $f^{abc}$ are the structure constants of the SU(3) Lie algebra. The quark field of flavor $f$ is denoted by $\psi_{f}$ and carries mass $m_{f}$. Its interaction with the gluon field is described by the gauge-covariant derivative $D_{\mu}=\partial_{\mu}-ig\,T^{a}A_{\mu}^{a},$ where $g$ is the strong coupling constant and $T^{a}$ are the generators of the SU(3) color group, satisfying the commutation relations $T^{a},T^{b}]=if^{abc}T^{c}.$
Because the QCD gauge field admits topologically nontrivial configurations, the most general gauge-invariant action allows a term proportional to $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}$. Its coefficient $\theta$ is the QCD vacuum angle and controls explicit $CP$ violation. When the quark mass matrix $M$ carries complex phases, an anomalous axial $U(1)_{A}$ rotation shifts these phases into the topological term. As a result, the physically observable $CP$-violating parameter is $\bar{\theta}=\theta-\arg\!\det(M),$ where $M$ denotes the full quark mass matrix. Since $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}$ is odd under $CP$, QCD is $CP$-symmetric only when $\bar{\theta}=0$, any nonzero value of $\bar{\theta}$ therefore leads to explicit $CP$ violation.
@@
In QCD, such a violation would be expected to generate low-energy hadronic effects. For example, it induces a permanent electric dipole moment (EDM) of the neutron [7, 8]. Precision measurements, however, have not observed such a signal. The current experimental bound places the neutron EDM $|d_{n}|<1.8\times 10^{-26}e\cdot\mathrm{cm}\ (90\%{\rm C.L.})$ [9, 10], which implies an extraordinarily small value of $|\bar{\theta}|\lesssim 10^{-10}$. This extreme suppression cannot be explained within the Standard Model, leading to the strong $CP$ problem: QCD allows a large $\bar{\theta}$, yet nature appears to choose a value extremely close to zero.
@@
A compelling theoretical resolution is provided by the Peccei-Quinn mechanism [11, 12, 13, 6], which promotes $\bar{\theta}$ to a dynamical field associated with a spontaneously broken global $U(1)_{\text{PQ}}$ symmetry. The resulting pseudo-Nambu-Goldstone boson, known as the axion, dynamically relaxes the effective $\bar{\theta}$ parameter to zero, thereby restoring $CP$ symmetry in the strong interactions [6]. The axion $a(x)$ couples anomalously to gluons
$\mathcal{L}_{a}=\frac{g^{2}}{32\pi^{2}}\frac{a}{f_{a}}G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu},$ (2)
where $f_{a}$ is the decay constant. As a result, the effective $CP$-violating angle becomes $\theta_{\rm eff}=\bar{\theta}+\frac{a}{f_{a}}$. Nonperturbative QCD dynamics generate an effective potential $V(\theta_{\rm eff})$ for this combination, which guarantee that the vacuum energy is minimized at $\theta_{\rm eff}=0$ independent of microscopic details. Consequently, the axion acquires a vacuum expectation value $\langle a\rangle=-\bar{\theta}f_{a}$ dynamically canceling the $CP$-violating phase. After shifting $a=a_{\mathrm{phys}}+\langle a\rangle$, the QCD vacuum is exactly $CP$-symmetric, while the physical axion $a_{\mathrm{phys}}$ remains as a light excitation around the minimum. In this way, the Peccei-Quinn mechanism does not rely on fine-tuning or an exact symmetry of the Standard Model. Instead, $CP$ conservation in the strong sector emerges dynamically from vacuum alignment, with the axion providing the necessary degree of freedom. The axion framework not only resolves the strong $CP$ problem but also provides a well-motivated candidate for dark matter, linking fundamental particle physics with cosmological observations [14].
@@
Theoretical progress on the strong-$CP$ problem is limited by the fact that the QCD vacuum is intrinsically nonperturbative and difficult to treat with conventional techniques. Conventional lattice methods struggle with real-time dynamics and $\bar{\theta}$-dependence, while analytical techniques rely on approximations. Quantum simulation [15, 16] provides a complementary route: instead of performing numerical sampling, one implements a controllable quantum system whose Hilbert space directly encodes the gauge theory and then measures how its vacuum responds to a $CP$-violating background. The (1+1)(1{+}1)-dimensional Schwinger model retains essential features such as confinement and topological vacuum sectors, yet is sufficiently simple to be mapped to a small number of qubits and executed on near-term hardware.
@@
In this work, we use the (1+1)(1{+}1)-dimensional Schwinger model to examine the microscopic origin of strong $CP$ violation and its axionic restoration. We construct a qubit-encoded Schwinger model with a $\bar{\theta}$-dependent electric sector, extend it to include a dynamical axion field, and prepare its vacuum using a feedback-driven quantum optimization algorithm (FALQON) [17]. By computing the vacuum energy as a function of $\bar{\theta}$, we first characterize the $\bar{\theta}$-dependent vacuum structure of the lattice gauge theory in the absence of an axion, including distortions arising from reduced dimensionality and finite-size effects. Upon coupling the system to a dynamical axion field, the vacuum energy relaxes to a minimum at $\theta_{\rm eff}=0$ and becomes nearly independent of the initial $\bar{\theta}$. These results demonstrate that the Peccei-Quinn mechanism can be realized in controlled, few-qubit quantum simulations, opening a new route to exploring topological gauge dynamics on quantum hardware.
@@
Gauge-field Lagrangian to Hamiltonian. To express the QCD Lagrangian in a form suitable for canonical quantization, we rewrite the gluon field-strength tensor $G_{\mu\nu}^{a}$ in terms of its temporal and spatial components. In analogy to electromagnetism, we define the chromoelectric and chromomagnetic fields as
@@
Substituting these definitions into the gauge-field part of the QCD Lagrangian (Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism), i.e., $\mathcal{L}_{\text{gauge}}=-\frac{1}{4}G_{\mu\nu}^{a}G^{a,\mu\nu}+\bar{\theta}\frac{g^{2}}{32\pi^{2}}G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu},$ and using the identity $G_{\mu\nu}^{a}\tilde{G}^{a,\mu\nu}=-4\mathbf{E}^{a}\cdot\mathbf{B}^{a}$, we obtain
$\mathcal{L}_{\text{gauge}}=\frac{1}{2}\mathbf{E}^{a}\cdot\mathbf{E}^{a}-\frac{1}{2}\mathbf{B}^{a}\cdot\mathbf{B}^{a}+\bar{\theta}\frac{g^{2}}{8\pi^{2}}\mathbf{E}^{a}\cdot\mathbf{B}^{a}.$ (4)
@@
$\Pi_{i}^{a}=\frac{\partial\mathcal{L}_{\text{gauge}}}{\partial(\partial_{0}A_{i}^{a})}=E_{i}^{a}+\bar{\theta}\frac{g^{2}}{8\pi^{2}}B_{i}^{a},$ (5)
showing that the $\bar{\theta}$-term shifts the momentum by an amount proportional to the chromomagnetic field. This term couples the electric and magnetic sectors, thus $\Pi_{i}^{a}$ no longer equals the electric field, and thereby modifying the canonical structure of the theory.
The Hamiltonian is then constructed via the Legendre transformation [18]
@@
$=\frac{1}{2}\left(\mathbf{E}^{a}\cdot\mathbf{E}^{a}+\mathbf{B}^{a}\cdot\mathbf{B}^{a}\right)-\bar{\theta}\frac{g^{2}}{8\pi^{2}}\mathbf{E}^{a}\cdot\mathbf{B}^{a}.$ (6)
@@
Consequently, the term proportional to $\bar{\theta}\mathbf{E}^{a}\cdot\mathbf{B}^{a}$ in the Hamiltonian (Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism) explicitly breaks $CP$ symmetry. Although small in magnitude (given experimental constraints on $\bar{\theta}$), this term could induce observable effects such as a permanent EDM in color-neutral hadrons like the neutron [7, 8].
Simplified (1+1)D U(1) Schwinger model. We consider a (1+1)-dimensional Schwinger model, corresponding to one spatial and one temporal dimension [19, 20]. Despite its Abelian gauge structure, this model reproduces several nonperturbative features in QCD, including confinement of charges, a dynamically generated mass gap, and spontaneous chiral symmetry breaking in the massless limit [20, 21, 22, 23]. It also possesses a nontrivial topological vacuum structure analogous to the $\bar{\theta}$-vacua of QCD, allowing controlled investigations of $CP$-violating effects induced by a background $\bar{\theta}$-term in (1+1) dimensions [23, 24]. Crucially, the Schwinger model admits a well-defined lattice discretization and a finite-dimensional gauge-link formulation [25], making it ideally suited for implementation on quantum computing platforms [26].
In (1+1) dimensions, the gauge field has only one spatial component, and the magnetic field identically vanishes. The gauge-field dynamics are therefore governed by the electric field $E(x)$, which serves as the canonical conjugate to the spatial gauge potential $A_{1}(x)$. The commutation relation reads $A_{1}(x),E(y)]=i\delta(x-y)$, analogous to that between position and momentum in quantum mechanics. The continuum Lagrangian for a single fermion flavor coupled to this Abelian gauge field, including the topological $\bar{\theta}$-term, can be written as
$\mathcal{L}=\bar{\psi}\big(i\gamma^{\mu}D_{\mu}-m\big)\psi-\frac{1}{2}E^{2}+\frac{\bar{\theta}}{2\pi}E,$ (9)
where the gauge-covariant derivative is $D_{\mu}=\partial_{\mu}-igA_{\mu}$. The Legendre transformation gives
$H=\int dx\Big[\psi^{\dagger}(x)\big(-i\gamma^{1}D_{1}+m\gamma^{0}\big)\psi(x)+\tfrac{1}{2}\Big(E(x)-\tfrac{\bar{\theta}}{2\pi}\Big)^{2}\Big],$
where the shift $E(x)\!\to\!E(x)-\tfrac{\bar{\theta}}{2\pi}$ encodes the $CP$-odd topological term. This structure is analogous to the $\mathbf{E}\!\cdot\!\mathbf{B}$ coupling in QCD but simpler in one spatial dimension.
For quantum simulation, we adopt the Kogut-Susskind lattice discretization [27, 28], with gauge links $U_{x,x+1}=e^{iA_{x,x+1}}$ and electric fields $E_{x}$. The discretized Hamiltonian reads
@@
$+\frac{g^{2}}{2}\sum_{x}\Big(E_{x}-\frac{\bar{\theta}}{2\pi}\Big)^{2},$ (10)
where $Z_{x}\equiv 1-2\psi_{x}^{\dagger}\psi_{x}$ denotes the local charge operator. The alternating sign $(-1)^{x}$ stems from the staggered-fermion scheme, which suppresses fermion doubling while preserving chiral symmetry.
@@
Two sites lattice case. Restricting the system to its smallest nontrivial instance, i.e., two sites lattice connected by a single link, yields the minimal Hamiltonian
@@
$+\dfrac{g^{2}}{2}\left(E-\frac{\bar{\theta}}{2\pi}\right)^{2}+\lambda\sum_{x}G_{x}^{2}.$ (12)
In this representation, the first term corresponds to the staggered mass energy, the second encodes gauge-invariant fermion hopping between the two sites mediated by the link operator $U_{01}$, and the third term represents the electric-field energy including the $CP$-violating $\bar{\theta}$ shift. The final term enforces Gauss’s law through an energetic constraint, restricting the evolution to the gauge-invariant subspace.
This minimal Schwinger model retains the topological and $CP$-violating structure of the QCD Hamiltonian. The $\bar{\theta}$-dependent electric-field shift plays the role of an effective $CP$-odd background field, allowing one to explore parity violation, vacuum degeneracy, and $\bar{\theta}$-vacuum transitions within a low-dimensional analog of strong interactions.
To implement the model on quantum computers, both the fermionic and gauge degrees of freedom are encoded in qubits (spin-$\tfrac{1}{2}$ systems). The fermionic modes are mapped to qubits using the Jordan-Wigner (JW) transformation [29], which preserves the correct anticommutation relations through parity strings, and each gauge link is represented by a single qubit in the quantum-link formulation. The staggered fermion fields $\psi_{x}$ and $\psi_{x}^{\dagger}$ satisfy
@@
where $S_{x}^{\pm}=(X_{x}\pm iY_{x})/2$. The string $\prod_{y<x}Z_{y}$ enforces fermionic parity between sites. Here $X$, $Y$, and $Z$ are Pauli matrices. The quantum-link representation of the gauge field is given by $E=(I-Z_{\ell})/2$ and $U=(X_{\ell}+iY_{\ell})/2$, where $\ell$ stands for the single-link qubit. See App. A for the full derivation.
@@
The electric-field energy term, including the $CP$-violating $\bar{\theta}$-shift yields
$H_{\bar{\theta}}$ $=\dfrac{g^{2}}{2}\Big(E-\dfrac{\bar{\theta}}{2\pi}\Big)^{2}$
$=\frac{g^{2}}{2}\left[\left(\frac{\bar{\theta}}{2\pi}\right)^{2}-\frac{\bar{\theta}}{2\pi}+\frac{1}{2}\right]I_{l}+\frac{g^{2}}{2}\left(\frac{\bar{\theta}}{2\pi}-\frac{1}{2}\right)Z_{\ell}.$ (16)
The constant term merely shifts the global energy, while the $Z_{\ell}$-dependent term changes sign at $\bar{\theta}=\pi$, leading to the correct $2\pi$ periodicity.
@@
Despite its simplicity, this model retains the nontrivial topological and $CP$-violating structure of QCD and provides an ideal platform for near-term quantum simulations of gauge-theory dynamics.